Using the CPA, the Deepwater Co. was able to determine that the expected completion
time for the construction of a racing yacht was 1 year 9 months and the project
variance was 4 months.
Find the probability that the project would be completed in
(a) 17 months
(b) 20months (hint: state your assumptions clearly
1
Expert's answer
2021-10-07T23:42:14-0400

Part a

"Z score =\t\n\\frac{x - \u03bc}{\n\u03c3}\\\\\n=\t\n\\frac{17 - 21}{\n2}\\\\\n=\t-2\\\\\nP-value \\space from \\space Z-Table:\\\\\n\nP(x<17) = 0.02275\\\\\n\nP(x>17) = 1 - P(x<17) = 0.97725\\\\\n\nP(17<x<21) = 0.5 - P(x<17) = 0.47725\\\\"


Part b

A probability distribution is used to define the potential values for each unknown variable in a simulation or assumption. The type of distribution you choose is determined by the variables surrounding it.

"Z score =\\frac{\t\nx - \u03bc}{\n\u03c3}\\\\\n=\\frac{\t\n20 - 21}{\n2}\\\\\n=\t-0.5\\\\\nP-value \\space from \\space Z-Table:\\\\\n\nP(x<20) = 0.30854\\\\\n\nP(x>20) = 1 - P(x<20) = 0.69146\\\\\n\nP(20<x<21) = 0.5 - P(x<20) = 0.19146"